# CAPTURe: Evaluating Spatial Reasoning in Vision Language Models via Occluded Object Counting

### v1

@@ C.2 Only Occluded Object Counting
Given that models seem to be able to identify patterns, we separately test whether models can count only the occluded objects in an image, which would be the next piece of reasoning needed to achieve accurate an accurate total. Here, as Tab. 7 demonstrates, the models perform especially poorly in this task, with high error rates across all models. Therefore, we can conclude that occlusion and counting are uniquely difficult for the VLMs, and that the drop in performance between unoccluded and occluded settings in Tab. 2 is more likely due to a poor ability to count occluded objects rather than to determine the pattern.
Model | Error (%) [] |
GPT-4o | 26.13 |
InternVL2 | 75.82 |
Molmo | 96.79 |
Qwen2VL | 32.89 |
Avg. of 4 VLMs | 57.91 |
C.3 Uncertainty/Confidence
Models are overconfident in occluded settings.
@@ C.3 Uncertainty/Confidence
We test the uncertainty with two different methods of obtaining confidence on Qwen2VL. In the first method, we prompt Qwen2VL for its confidence in the answer. For the second method, we generate 20 responses for every question in our VQA and calculate the confidence as the percentage of times the most common answer was generated. These results can be seen in Fig. 10 and Fig. 9 respectively. In both reliability curves, there is a slight trend that the models confidence is negatively correlated with the error, which is the desired outcome. In CAPTURe, however, the correlation is much stronger. While the models are somewhat calibrated (with generally lower confidence on higher-error examples, there are still outliers in prompted confidence for CAPTURe occluded and sampled confidence for CAPTURe occluded. This indicates that not only do the models perform worse under occlusion, they can also be overconfident.
C.4 Confusion Matrix: Predicted vs. Ground Truth Counts
We generate confusion matrices for every model on CAPTURe occluded images to investigate whether models are biased to predict certain numbers. The y-axis represents the ground truth values and the x-axis represents the model’s answers. Fig. 11 demonstrates that GPT-4o, InternVL2, and Qwen2VL are more likely to predict numbers that occur more frequently in text.
@@ Appendix D VLM Prompts
For each setting, we use a 100-example validation set to select the best prompt, which we report below.

### v2

@@ C.3 Models struggle at counting just occluded objects
We separately test whether models can count only the occluded objects (not including the visible objects) in an image. Here, as Tab. 8 demonstrates, the models perform especially poorly in this task, with high error rates across all models. Therefore, we can conclude that occlusion and counting are uniquely difficult for the VLMs, and that the drop in performance between unoccluded and occluded settings in Tab. 2 is likely due to a poor ability to count occluded objects.
Model | Error (%) [] | |
All Objects | Only Occluded | |
GPT-4o | 14.75 | 26.13 (+11.38) |
InternVL2 | 32.90 | 75.82 (+42.92) |
Molmo | 32.49 | 96.79 (+64.30) |
Qwen2VL | 29.33 | 32.89 (+3.56) |
Avg. of 4 VLMs | 27.37 | 57.91 (+30.54) |
C.4 Models are overconfident in occluded settings
@@
We test the uncertainty with two different methods of obtaining confidence on Qwen2VL. In the first method, we prompt Qwen2VL for its confidence in the answer. For the second method, we generate 20 responses for every question in our VQA and calculate the confidence as the percentage of times the most common answer was generated. These results can be seen in Fig. 9 and Fig. 10 respectively. In both reliability curves, there is a slight trend that the model’s confidence is negatively correlated with the error, which is the desired outcome. In CAPTURe, however, the correlation is much stronger. While the models are somewhat calibrated (with generally lower confidence on higher-error examples, there are still outliers in prompted confidence for CAPTURe occluded and sampled confidence for CAPTURe occluded. This indicates that not only do the models perform worse under occlusion, but they can also be overconfident.
C.5 Models are biased to predict specific numbers.
To examine where models frequently err, we generated a confusion matrix for every model based on CAPTURe results (shown in Sec. C.5). The y-axis represents the ground truth values and the x-axis represents the model’s answers. We find that models often over-predict numbers associated with common counts in real life: GPT-4o tends to predict numbers like 8, 9, 10, and 12, which are all non-prime numbers (i.e. can be arranged into a grid) and common groupings of objects. For example, 12 is a common grouping (dozens) and allows arrangements into 3x4 or 2x6 grids. InternVL and Qwen2VL over-predict 5 and 10, aligning with how humans conceptualize numbers. Indeed, Coupland [12] found that numbers 5, 10, 20, and other round numbers appear disproportionally more in online texts. Molmo has no correlation with these factors, possibly due to its unique “point and count” ability.
@@ Appendix D VLM Prompts
We use a 100-example validation set for each setting to select the best prompt, which we report below.